# Metrics on completely positive maps via noncommutative geometry

### v1

@@ Abstract
By considering an infinite-dimensional analogue of the Choi-Jamiolkowski isomorphism, we study how to induce metrics on a distinguished subset of the completely positive maps between tracial $C^{*}$-algebras using seminorms from noncommutative geometry. Under suitable conditions on the these seminorms, we show that the induced metrics will satisfy the quantum information theoretic properties of stability and chaining. Lastly, we show how to generate such metrics from Kasparov exterior products of spectral triples.
footnotetext: Emails: areaus@math.uio.no, bedos@math.uio.no, jonaeid@math.uio.no, nadiasl@math.uio.no, and trono@math.uio.no.
1 Introduction
@@
Trace-preserving completely positive maps are ubiquitous in Quantum Information Theory as, among others, the means to model evolution of quantum systems, these generally being described by matrix algebras. Through a standard duality perspective, there is a companion class of unital, completely positive maps, and these in turn are prominent in Operator Algebras and Noncommutative Geometry, especially in relation to deep questions about approximation properties of operator algebras (in an infinite-dimensional setting). Among many questions which have been asked about trace-preserving completely positive maps one finds those about existence of a metric structure, or distance, capable of capturing permanence properties such as chaining, a form of coherent composition, or stability under tensorial expansion of the system with an ancilla system [BV25, GLN05] present. In this work we propose a framework for developing a metric structure on the set of (unital) completely positive maps between two, not necessarily finite-dimensional $C^{*}$-algebras, by employing constructions native to Noncommutative Geometry.
@@
Attempts so far at developing a metric structure have concentrated on the finite-dimensional setting, namely matrix algebras [BV25], or, in the operator algebraic framework, on the von Neumann algebraic setting, where normal states are a core ingredient [AJL25, AJW25, Daw25].
The main thrust of the present work is that by employing completely positive maps between tracial $C^{*}$-algebras and formalizing a $C^{*}$-algebraic version of an embedding akin to the Choi-Jamiolkowski map in the finite-dimensional setting, we are in a position to take advantage of the constructions of quantum metrics, or distances, on state spaces of $C^{*}$-algebras pioneered by Connes and Rieffel, see e.g.[Con89, Rie98, Rie99, Rie02, Rie04].
Explicitly, we specify a suitable subset of the completely positive maps between $C^{*}$-algebras $A$ and $B$, where $A$ is unital and $B$ is endowed with a faithful trace $\tau_{B}$, which we call trace channels (see Definition 4.13). These are exactly the completely positive maps which yield states under the infinite-dimensional version of the Choi-Jamiolkowski map. We are then able to construct metrics on the space of trace channels by employing techniques from noncommutative geometry, and put forward definitions of stability and chaining in an infinite-dimensional setting. Moreover, we find sufficient conditions for stability and chaining to hold, and illustrate their validity in naturally occuring examples by once more appealing to constructions in noncommutative geometry. Indeed, we show stability for quantum metrics arising from spectral triples (cf. Theorem 6.2), and prove that chaining is satisfied for a natural class of maps between group $C^{*}$-algebras equipped with metrics arising from spectral triples coming from proper length functions on the group, see Theorem 6.9. A large class of examples exhibiting the best possible behavior where chaining and stability hold simultaneously is provided by countable amenable groups with natural length functions, see Example 6.10. The construction of Lipschitz seminorms coming from spectral triples is of importance in the study of compact quantum metric spaces in noncommutative geometry. It has been investigated in a variety of different settings, e.g. for group $C^{*}$-algebras [CR17, OR05], crossed products [AKK25, Haw+13, Kli24], and quantum groups [AK18, AK26, BVZ15, KK25], though this is far from a complete list. In a related line of research, [AJL25] investigates Lipschitz seminorms (which a priori do not come from spectral triples) on quantum permutation groups for the purpose of studying the quantum Wasserstein distance.
Questions such as having stability and chaining are not standard in the context of quantum metrics in noncommutative geometry, and we hope this work opens up a new fruitful line of interaction with quantum information theory.
@@
As mentioned, a large part of the present work relies on formalizing an infinite-dimensional analogue of the Choi-Jamiolkowski map for $C^{*}$-algebras. Related work has appeared in several recent papers: In the quantum graph literature an analogue appears for finite-dimensional $C^{*}$-algebras, cf. [Was24, CGW25]. In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiolkowski isomorphism, see [Haa21, HKS24] and references therein.
@@
The technical difficulties one encounters when moving from a von Neumann algebraic setup to a $C^{*}$-algebraic setting pertain mainly to the need to work with state spaces of tensor products of $C^{*}$-algebras and their opposite algebras. This difficulty does not appear in the finite-dimensional setting as $M_{n}(\mathbb{C})$ is nuclear and isomorphic to its opposite algebra. The latter of these facts is not true for general $C^{*}$-algebras. Thus, without requiring nuclearity of the $C^{*}$-algebras in question, certain maps and their continuity properties very much depend on whether one employs the minimal or the maximal tensor product. A celebrated result of Kirchberg, cf. [Kir93, Kir94], generalizing a landmark result of Connes, cf. [Con76], establishes continuity with respect to the minimal tensor product of a natural map $\mu_{\tau}$, cf. Theorem 2.2. The map $\mu_{\tau}$ is dependent on a trace $\tau$, and Kirchberg’s Theorem states that $\mu_{\tau}$ is continuous in the spatial norm if and only if the corresponding trace $\tau$ is amenable. This allows us to define a tractable version of the Choi-Jamiolkowski map for maps between $C^{*}$-algebras, where the target $C^{*}$-algebra is unital and admits an amenable trace, cf. Theorem 4.15.
@@
The structure of the paper is as follows. We collect a number of important notions and set notation in the preliminary section Section 2. While we focus on a specific way of inducing metrics on certain subsets of completely positive maps in this article, this is not the only induction scheme one could employ. As such, we discuss induction of metrics in Section 3. It also serves to explain the motivation behind the remaining sections. In Section 4 we study the $C^{*}$-algebraic Choi-Jamiolkowski map and use it to introduce the notion of trace channels. The $C^{*}$-algebraic Choi-Jamiolkowski map is then used in Section 5 to induce metrics on the set of trace channels between $C^{*}$-algebras. We especially focus on criteria guaranteeing that these metrics satisfy stability and chaining. Lastly, in Section 6 we use machinery from noncommutative geometry to show that the necessary conditions for stability and chaining from Section 5 are automatically satisfied for metrics arising from the external Kasparov product of spectral triples. We in particular focus on spectral triples for group $C^{*}$-algebras arising from proper length functions on countable groups.
@@ Acknowledgments
This research was funded by The Research Council of Norway [project 324944] and [project 345433].
2 Preliminaries
2.1 Tensor norms, opposite algebras, and traces
@@
If $A$ and $B$ are $C^{*}$-algebras, then the algebraic tensor product $A\odot B$ of $A$ and $B$ becomes a *-algebra with respect to the involution determined by $(a\otimes b)^{*}=a^{*}\otimes b^{*}$ for all $a\in A$ and $b\in B$ (see for instance [Mur90], p. 188-189). Moreover, if $\delta$ is a $C^{*}$-norm on $A\odot B$, then we will denote the completion of $A\odot B$ with respect to $\delta$ by $A\otimes_{\delta}B$. There are two $C^{*}$-norms in particular that are natural to consider.
@@
where $\pi_{A}\colon A\to B(H_{A})$ is any faithful *-representation of $A$, and $\pi_{B}\colon B\to B(H_{B})$ is any faithful *-representation of $B$. It is well-known that the spatial norm is independent of the choices of $\pi_{A}$ and $\pi_{B}$. The spatial norm satisfies $\|x\|_{\rm min}\leq\delta(x)$ for any $x\in A\odot B$ and any $C^{*}$-norm $\delta$ on $A\odot B$. This is known as Takesaki’s Theorem cf. [Mur90, Section 6.4] or [BO08, Section 3.4] for a proof. We denote the completion of $A\odot B$ with respect to the spatial norm by $A\otimes_{\rm min}B$ and call this the minimal tensor product of $A$ and $B$.
@@
By a trace $\tau$ on a $C^{*}$-algebra $B$ we will mean that $\tau$ is a nonzero tracial positive functional on $B$. In the event that $\tau$ is normalized, hence a state on $B$, we will call $\tau$ a tracial state. We make this distinction to ease the exposition, and in the case where we discuss matrix algebras, $M_{n}(\mathbb{C})$, the most natural trace to consider is the unnormalized one.
If $B$ is a $C^{*}$-algebra, we will denote by $B^{\textup{op}}=\{b^{\rm op}:b\in B\}$ the opposite $C^{*}$-algebra of $B$, with product given by $b^{\rm op}c^{\rm op}=(cb)^{\rm op}$ for all $b,c\in B$. The multiplication in $B$ induces a linear map $\mu\colon B\odot B^{\rm op}\to B$ defined by
@@
It is natural to ask for which $C^{*}$-norms on $B\odot B^{\rm op}$ the map $\mu_{\tau}$ is continuous. For the maximal norm we have the following result which follows by elementary methods, cf. [Oza04, Page 21] for an exposition.
@@
This is only two of several equivalent statements proved by Kirchberg distributed over the papers [Kir93] and [Kir94]. A proof of these equivalences can be found in [BO08, Theorem 6.2.7]. Note that by amenability of $\tau$ we mean that the tracial state $\frac{1}{\tau(1_{B})}\tau$ is amenable in the following sense:
@@
Whenever $\mu_{\tau}$ is continuous with respect to some $C^{*}$-norm $\delta$, we will use the notation $\mu_{\tau}^{\delta}$ to denote the (unique) extension of $\mu_{\tau}$ to $B\otimes_{\delta}B^{\rm op}$. In the case where $B$ is nuclear, no extra decoration will be made to denote the extension of $\mu_{\tau}$ to $B\otimes B^{\rm op}$. In the case where $\delta=\|\cdot\|_{\rm min}$ or $\delta=\|\cdot\|_{\rm max}$, we will use the notation $\mu_{\tau}^{\rm min}$ or $\mu_{\tau}^{\rm max}$ respectively.
@@
extends to a completely positive map on both the minimal and maximal tensor products. Denoting (either) of the extensions by $F\otimes G$ we have have that
@@
for any $\psi\in B^{*}$.
For a unital $C^{*}$-algebra $A$, we will denote the unit by $1_{A}$. In the special case where $A=M_{n}(\mathbb{C})$, we will for simplification use the notation $1_{n}$ instead of $1_{M_{n}(\mathbb{C})}$. Similarly, we will use the notation $\operatorname{id}_{n}$ for the identity map on $M_{n}(\mathbb{C})$.
2.2 Lipschitz seminorms and spectral triples
We remind the reader about basic notions from quantum metric geometry, in particular the definition of a Lipschitz seminorm. There are numerous definitions in the literature, e.g. [Rie04, Rie04a, Rie16, Kaa24], all of which are related and serve different purposes. In the sequel we will find it beneficial to use the definition from [KK25].

### v2

@@ Abstract.
We study methods of inducing metrics on unital completely positive maps by employing seminorms arising in noncommutative geometry. Our main approach relies on the development of an infinite-dimensional $C^{*}$-algebraic analogue of the Choi-Jamiołkowski isomorphism. Under suitable conditions, we show that the induced metrics satisfy the quantum information theoretic properties of stability and chaining. Moreover, we show how to generate such metrics using constructions native to noncommutative geometry, by for example using external Kasparov products of spectral triples.
Key words and phrases:
Stability, chaining, quantum metrics, unital completely positive maps, spectral triples, external Kasparov products, twisted group C*-algebras, quantum information theory, noncommutative geometry
2020 Mathematics Subject Classification:
Primary: 81R60 ; Secondary: 58B34, 81P45, 81P47, 46L89
1. Introduction
@@
Trace-preserving completely positive maps are ubiquitous in quantum information theory as, among others, the means to model evolution of quantum systems, these generally being described by matrix algebras. Through a standard duality perspective, there is a companion class of unital, completely positive maps, and these in turn are prominent in operator algebras and noncommutative geometry, especially in relation to deep questions about approximation properties of operator algebras (in an infinite-dimensional setting). Among many questions which have been asked about trace-preserving completely positive maps one finds those about existence of a metric structure, or distance, capable of capturing permanence properties such as chaining, a form of coherent composition, or stability under tensorial expansion of the system with an ancilla system present, cf. [GLN05, BV25]. In this work we propose a framework for developing a metric structure on the set of (unital) completely positive maps between two, not necessarily finite-dimensional $C^{*}$-algebras, by employing constructions native to noncommutative geometry.
@@
Attempts to develop a metric structure have so far concentrated on the finite-dimensional setting, namely matrix algebras [GLN05, BV25], or, in the operator algebraic framework, on the von Neumann algebraic setting, where normal states are a core ingredient [AJL25, AJW25, Daw25]. An exception to this is the Wasserstein type metric of order one on unital completely positive maps between tensor products of unital $C^{*}$-algebras introduced in [DM23].
The main thrust of the present work is the formalization and study of two methods of inducing metrics on the set of unital completely positive maps between $C^{*}$-algebras through the use of seminorms, providing a rich source of examples coming from noncommutative geometry. In particular, it provides a link to the theory of compact quantum metric spaces, as pioneered by Connes and Rieffel, see, e.g., [Con89, Rie98, Rie99, Rie02, Rie04]. Very satisfyingly, we find that both stability and chaining have natural and precise formulations and, moreover, are in place for large classes of examples. Questions regarding metrics having stability and chaining are not standard in the context of quantum metrics in noncommutative geometry, and we hope that this work opens up a new fruitful line of interaction with quantum information theory.
The investigation of these questions in such a large generality necessitates the development of novel technical tools not previously found in the literature. Of note, we develop a $C^{*}$-algebraic Choi-Jamiołkowski type embedding to account for the fact that unlike matrix algebras, general $C^{*}$-algebras are neither nuclear nor necessarily isomorphic to their opposite algebras.
Before we delve into the details of the paper, we take a moment to briefly describe the origins of the notions of stability and chaining. In [GLN05], see also [BV25], the authors outline a program to define distances, or metrics, on sets of quantum processes, in particular covering quantum channels, i.e., trace-preserving completely positive maps between matrix algebras $A$ and $B$. Several potentially meaningful distances are put forward, along with a list of desirable properties these distances should satisfy. The first of two particularly important properties a (possibly extended) metric $d$ should satisfy is stability, in the sense that if $\mathcal{F},\mathcal{G}$ are quantum channels from $A$ to $B$, then $d(\operatorname{id}_{n}\otimes\mathcal{F},\operatorname{id}_{n}\otimes\mathcal{G})=d(\mathcal{F},\mathcal{G})$ must hold for every $n\in\mathbb{N}$. Here $\operatorname{id}_{n}\otimes\mathcal{F}$ and $\operatorname{id}_{n}\otimes\mathcal{G}$ are amplifications of the initial channels acting between $M_{n}(\mathbb{C})\otimes A$ and $M_{n}(\mathbb{C})\otimes B$. From a quantum information theoretic perspective, this property is supposed to reflect the fact that an unrelated ancillary system should not affect the distance between initial quantum channels modelling quantum processes. In [GLN05] and [BV25] it is expected that an analogue of $d$ exists on all amplifications for every $n\in\mathbb{N}$.
The second key property a distance on quantum channels should feature is chaining, which aims to capture the idea that if one divides a process into smaller steps, the total error should be smaller than the sum of the stepwise errors. Here one interprets the distance between quantum processes as a measure of the error between an ideal process and an observed process. Roughly, this is expressed as $d(\mathcal{G}_{1}\circ\mathcal{F}_{1},\mathcal{G}_{2}\circ\mathcal{F}_{2})\leq d(\mathcal{G}_{1},\mathcal{G}_{2})+d(\mathcal{F}_{1},\mathcal{F}_{2})$, assuming the quantities are meaningfully defined (see definition 3.2).
Among the metrics highlighted in [GLN05] as the most promising, especially with regard to stability and chaining, there are one arising from a stabilizing procedure and one based on the Choi-Jamiołkowski isomorphism. The more recent overview paper [BV25] reinforces this view, while also presenting some new definitions of distances.
Our concern here is with metrics on sets of completely positive maps. The set $\operatorname{UCP}(A,B)$ of unital completely positive maps between unital $C^{*}$-algebras $A$ and $B$ is known to be dual to the class of quantum channels from $B$ to $A$ (see for instance [JKP09, Wat09]). Thus we translate the discussions in [GLN05] and [BV25] to the setting of $\operatorname{UCP}(A,B)$ whenever relevant and without further mention. In this spirit, we put forward definitions of stability and chaining in an infinite-dimensional setting in section 3. We will establish these properties in two distinct directions: one involves a stabilizing process similar to [GLN05, BV25], but with the difference that it starts from an extended metric $d_{L}$ arising from a seminorm $L$ on a unital $C^{*}$-algebra $A$, see pullback-induced extended metrics in definition 3.3 and section 4. We are able to find sufficient conditions under which stability and chaining are satisfied for these extended metrics in theorem 4.5 and corollary 4.7, respectively.
Another way of inducing a metric on $\operatorname{UCP}(A,B)$ motivated by [GLN05] involves embedding $\operatorname{UCP}(A,B)$ into the state space of a higher-dimensional $C^{*}$-algebra using the Choi-Jamiołkowski isomorphism. If one naively assumes that such an embedding exists in general, say $\omega\colon\operatorname{UCP}(A,B)\to S(A\otimes_{\rm max}B^{\rm op})$ is such an embedding (in analogy with the constructions from [GLN05, BV25]), then, given a seminorm $L$ on $A\otimes_{\rm max}B^{\rm op}$, one can define an embedding-induced extended metric
$\Delta_{L}(F,G):=\operatorname{mk}_{L}(\omega(F),\omega(G))\text{ for }F,G\in\operatorname{UCP}(A,B),$
where $\operatorname{mk}_{L}$ is the Monge-Kantorovič extended metric associated with $L$. Monge-Kantorovič metrics are central to noncommutative geometry, see section 2.2 for a short introduction. However, such a Choi-Jamiołkowski embedding is not known to exist for general $C^{*}$-algebras. We develop such an embedding in section 5 where we consider a $C^{*}$-algebra $B$ with a faithful trace $\tau$, and we define the map $\omega_{\tau}\colon\operatorname{Hom}(A,B)\to\operatorname{Hom}(A\odot B^{\rm op},\mathbb{C})$ which gives an infinite-dimensional analogue of the map appearing in the Choi-Jamiołkowski isomorphism. This map always extends to give a well-defined map into $\operatorname{Hom}(A\otimes_{\rm max}B^{\rm op},\mathbb{C})$, under which $\operatorname{UCP}(A,B)$ embeds into $S(A\otimes_{\rm max}B^{\rm op})$, cf. theorem 5.15. In fact, theorem 5.15 precisely characterizes which maps from $A$ to $B$ that correspond to states on $A\otimes_{\rm max}B^{\rm op}$. We call these maps trace channels, see definition 5.13, and they will be the main object of study in section 5 and section 6.
@@
A large part of the present work consists of formalizing an infinite-dimensional analogue of the Choi-Jamiołkowski map for $C^{*}$-algebras. Related work has appeared in several recent papers: In the quantum graph literature an analogue appears for finite-dimensional $C^{*}$-algebras, cf. [Was24, CGW25]. In a von Neumann algebraic setting, there exist some infinite-dimensional versions of the Choi-Jamiołkowski isomorphism, see [Haa21, HKS24] and references therein.
@@
The technical difficulty one encounters when moving from a von Neumann algebraic setup to a $C^{*}$-algebraic setting pertains mainly to the need to work with state spaces of tensor products of $C^{*}$-algebras and their opposite algebras. This difficulty does not appear in the finite-dimensional setting as $M_{n}(\mathbb{C})$ is nuclear and isomorphic to its opposite algebra. The latter of these facts is not true for general $C^{*}$-algebras. Thus, without requiring nuclearity of the $C^{*}$-algebras in question, certain maps and their continuity properties very much depend on whether one employs the minimal or the maximal tensor product. A celebrated result of Kirchberg, cf. [Kir93, Kir94], generalizing a landmark result of Connes, cf. [Con76], establishes continuity with respect to the minimal tensor product of a natural map $\mu_{\tau}$, cf. theorem 2.2. The map $\mu_{\tau}$ is dependent on a trace $\tau$, and Kirchberg’s Theorem states that $\mu_{\tau}$ is continuous in the spatial norm if and only if the corresponding trace $\tau$ is amenable. This allows us to define a tractable version of the Choi-Jamiołkowski map for maps between $C^{*}$-algebras, where the target $C^{*}$-algebra is unital and admits an amenable trace, cf. theorem 5.15.
We have chosen to formulate our definition of a Choi-Jamiołkowski map for $C^{*}$-algebras in the most general case available to us although admittedly the scenario of non-nuclear $C^{*}$-algebras is perhaps of little appeal, or apparent relevance, to questions in quantum information theory. Nevertheless, we have done so in anticipation of further interaction between the fields.
We are then able to construct metrics on the space of trace channels by employing techniques from noncommutative geometry. Moreover, we find sufficient conditions for stability and chaining to hold, and illustrate their validity in naturally occurring examples by once more appealing to constructions in noncommutative geometry. Indeed, we show stability for extended metrics on trace channels arising from external Kasparov products of spectral triples (cf. theorem 7.2). Moreover, we prove that chaining is satisfied for extended metrics arising from spectral triples coming from proper length functions on countable amenable groups if we restrict the unital completely positive maps in question to those arising as Fourier multipliers. A large class of examples exhibiting the best possible behavior where chaining and stability hold simultaneously is provided by countable amenable groups with natural length functions, see example 7.8. The construction of seminorms coming from spectral triples is of importance in the study of compact quantum metric spaces in noncommutative geometry. It has been investigated in a variety of different settings, e.g. for group $C^{*}$-algebras [OR05, CR17], crossed products [Haw+13, Kli24, AKK25], and quantum groups [BVZ15, AK18, KK25, AK26], though this is far from a complete list. In a related line of research, [AJL25] investigates seminorms (which a priori do not come from spectral triples) on quantum permutation groups for the purpose of studying the quantum Wasserstein distance.
@@
The structure of the paper is as follows. We collect a number of important notions and set notation in section 2. In section 3 we formalize the notions of stability and chaining mathematically and discuss pullback-induced and embedding-induced extended metrics on sets of unital completely positive maps. We study stability and chaining for pullback-induced extended metrics in section 4. In section 5 we study the $C^{*}$-algebraic Choi-Jamiołkowski map and use it to introduce the notion of trace channels. This map is then used in section 6 to define and study embedding-induced extended metrics on the set of trace channels between $C^{*}$-algebras. We especially focus on criteria guaranteeing that these metrics satisfy stability and chaining. Lastly, in section 7 we use machinery from noncommutative geometry to show that the necessary conditions for stability and chaining from section 6 are automatically satisfied for metrics arising from the external Kasparov product of spectral triples. We in particular focus on spectral triples for twisted group $C^{*}$-algebras arising from proper length functions on countable amenable groups.
@@ Acknowledgments
This research was funded by The Research Council of Norway [project 324944] and [project 345433]. The third author acknowledges the financial support of The Norwegian National Security Authority.
2. Preliminaries
2.1. Tensor norms, opposite algebras, and traces
@@
If $A$ and $B$ are $C^{*}$-algebras, then the algebraic tensor product $A\odot B$ of $A$ and $B$ becomes a *-algebra with respect to the involution determined by $(a\otimes b)^{*}=a^{*}\otimes b^{*}$ for all $a\in A$ and $b\in B$ (see for instance [Mur90, p. 188-189]). Moreover, if $\delta$ is a $C^{*}$-norm on $A\odot B$, then we will denote the completion of $A\odot B$ with respect to $\delta$ by $A\otimes_{\delta}B$. There are two $C^{*}$-norms in particular that are natural to consider.
@@
where $\pi_{A}\colon A\to B(H_{A})$ is any faithful *-representation of $A$, and $\pi_{B}\colon B\to B(H_{B})$ is any faithful *-representation of $B$. It is well-known that the spatial norm is independent of the choices of $\pi_{A}$ and $\pi_{B}$. The spatial norm satisfies $\|x\|_{\rm min}\leq\delta(x)$ for any $x\in A\odot B$ and any $C^{*}$-norm $\delta$ on $A\odot B$. This is known as Takesaki’s Theorem, cf. [Mur90, Section 6.4] or [BO08, Section 3.4] for a proof. We denote the completion of $A\odot B$ with respect to the spatial norm by $A\otimes_{\rm min}B$ and call this the minimal tensor product of $A$ and $B$.
@@
By a trace $\tau$ on a $C^{*}$-algebra $B$ we will mean that $\tau$ is a nonzero tracial positive functional on $B$. When $\tau$ is normalized, hence a state on $B$, we will call $\tau$ a tracial state. We make this distinction to ease the exposition, and in the case where we discuss matrix algebras, $M_{n}(\mathbb{C})$, the most natural trace to consider is the canonical (unnormalized) one.
If $B$ is a $C^{*}$-algebra, we will denote by $B^{\textup{op}}=\{b^{\rm op}:b\in B\}$ the opposite $C^{*}$-algebra of $B$, with product given by $b_{1}^{\rm op}b_{2}^{\rm op}=(b_{2}b_{1})^{\rm op}$ for all $b_{1},b_{2}\in B$. The multiplication in $B$ induces a linear map $\mu\colon B\odot B^{\rm op}\to B$ defined by
@@
It is natural to ask for which $C^{*}$-norms on $B\odot B^{\rm op}$ the map $\mu_{\tau}$ is continuous. For the maximal norm we have the following result which follows by elementary methods, cf. [Oza04, p. 21] for an exposition.
@@
These are only two of several equivalent statements proved by Kirchberg distributed over the papers [Kir93] and [Kir94]. A proof of these equivalences can be found in [BO08, Theorem 6.2.7]. Note that by amenability of $\tau$ we mean that the tracial state $\frac{1}{\tau(1_{B})}\tau$ is amenable in the following sense:
@@
In general, we will denote the unique extension of $\mu_{\tau}$ to $B\otimes_{\rm max}B^{\rm op}$ by $\mu_{\tau}^{\rm max}$. Likewise, if $\mu_{\tau}$ extends to a continuous map on $B\otimes_{\rm min}B^{\rm op}$, we denote the extension by $\mu_{\tau}^{\rm min}$. In case $B$ is nuclear, no extra decoration will be made to denote the extension of $\mu_{\tau}$ to $B\otimes B^{\rm op}$.
@@
extends to a completely positive map on both the minimal and maximal tensor products. Denoting (either of) the extensions by $F\otimes G$ we have that
@@
for any $\psi\in B^{*}$. The state space of $A$ is denoted by $S(A)$.
For a unital $C^{*}$-algebra $A$, we will denote the unit by $1_{A}$. In the special case where $A=M_{n}(\mathbb{C})$, we will for simplification use the notation $1_{n}$ instead of $1_{M_{n}(\mathbb{C})}$. Similarly, we will use the notation $\operatorname{id}_{n}$ for the identity map on $M_{n}(\mathbb{C})$ and $\operatorname{Tr}_{n}$ for the (unnormalized) trace on $M_{n}(\mathbb{C})$. We furthermore use the notation $\left\{e_{i,j}^{(n)}\right\}_{i,j=1}^{n}$ to denote the standard matrix units of $M_{n}(\mathbb{C})$.
2.2. Lipschitz seminorms and spectral triples
We remind the reader about basic notions from quantum metric geometry, in particular the definition of a Lipschitz seminorm and of a compact quantum metric space. There are numerous definitions in the literature, e.g., [Rie04, Rie04a, Rie16, Kaa24], all of which are related and serve different purposes. In the sequel we will find it beneficial to use the terminology from [KK25].